Fred and George value fireworks differently. Fred’s demand for fireworks is Q = 100 − 5P, and George’s demand is Q = 90 − 3P. a] If the marginal cost of fireworks is $5, what is the socially optimal number of fireworks? b] Under Lindahl pricing, what share of the tax burden would Fred pay? What share would George pay
Answer #1

A) Freds Demand Q=100-5P. P=20-0.2Q

Geroges demand Q=90-3P. P=30-0.333Q

market demand for the public good is the vertical summation of individual demand curve.

P=50-0.5333Q

For socially optimal quantity P=MC

50-0.5333Q=5

Quantity=45/0.5333=84.427 and

B)Price of fred=20-0.2(84.427)=3.11444

price of George=30-0.333(84.427)=1.885

George share=1.885/5*100=37.716%